# Finch members

### v2

@@ Abstract
We introduce a finance & accounting benchmark (Finch) for evaluating AI agents on real-world, enterprise-grade professional workflows—interleaving data entry, structuring, formatting, web search, cross-file retrieval, calculation, modeling, validation, translation, visualization, and reporting. Finch is sourced from authentic enterprise workspaces at Enron (15,000 spreadsheets and 500,000 emails from 150 employees) and other financial institutions, preserving in-the-wild messiness across multimodal artifacts (text, tables, formulas, charts, code, and images) and spanning diverse domains such as budgeting, trading, and asset management.
@@
We conduct both human and automated evaluations of frontier AI systems including GPT 5.1, Claude Sonnet 4.5, Gemini 3 Pro, Grok 4, and Qwen 3 Max, and GPT 5.1 Pro spends 48 hours in total yet passes only 38.4% of workflows, while Claude Sonnet 4.5 passes just 25.0%. Comprehensive case studies further surface the challenges that real-world enterprise workflows pose for AI agents.
@@ 1 Introduction
However, real-world F&A work is inherently messy with substantial contextual complexity: artifacts are interconnected across heterogeneous spreadsheets, PDFs, and more, evolving through multiple versions with collaborative edits [klimt2004enron]; spreadsheets contain large, complex structures [spreadsheetllm2024] with cross-sheet references, intricate layouts, inconsistent formatting, cryptic terms, erroneous formulas, and multimodal artifacts such as charts, images, and code. It is also long-horizon [gdpval2025]: workflows demand multi-step reasoning spanning data entry, editing, retrieval, calculation, modeling, validation, reporting, and more.
@@
  • In-the-wild enterprise sourcing: Finch is built around authentic enterprise spreadsheets, emails, and PDFs from real-world enterprise workspaces—primarily Enron [enrondata_pst] (about 15,000 spreadsheet files and 500,000 emails from 150 executives and employees) and EUSES [fisher2005euses] (about 450 financial spreadsheets from various sources), along with securities and asset management firms, global organizations such as World Bank [worldbank_idr_2024], and Canadian and British governments [financecanada_fiscalref_2025, hmt_pesa_2023]. Documents are large, cross-referenced, and messy—containing rich multimodal artifacts such as tables, formulas, charts, pivots, and images.
  • Rigorous construction process: We propose a novel workflow construction pipeline grounded in real collaborative context of emails and versioned artifacts. We induce workflows from enterprise email threads and attachments, where collaborators naturally describe, discuss, and track workflows as part of their daily work. And we propose an LLM-assisted, expert-verified method to derive workflows by analyzing changes across versioned spreadsheets, surfacing the underlying goals that drive professionals’ work. Annotators must reason over large multi-sheet workbooks, and subtle version deltas to infer what the original analyst was trying to achieve, making the annotation process substantially more difficult than curating QA pairs over isolated tables.
@@
Figure 3: Illustration of an end-to-end predictive modeling workflow commonly performed by financial analysts. The workflow involves multiple steps, including web search, data import, cross-sheet and cross-file retrieval, calculation and financial modeling, visualization, and report generation. More illustrative examples for data characteristics in Finch are presented in Appendix E.
@@ 2.1 Dataset Construction
All annotated workflows from these different channels are consolidated into a unified schema with consistent fields (NL instruction, input files, reference outputs), and each workflow is tagged with task types (e.g., data entry/import, structuring, validation) and business types (e.g., planning and budgeting, pricing and valuation, operations, asset management). Note that the reference outputs include both file-based reference answers (for most generation/editing cases) and textual reference answers (for a small number of QA and summary/visualization cases).
@@ 2.1.1 Workflow from Enterprise Email Threads
We first mine enterprise email threads to surface workflows. Starting from the Enron Email Corpus, we prompt GPT-5 to identify collaborative messages that (i) explicitly state a business goal (e.g., “update the RAC rankings” or “revise the 2002 allocations”) and (ii) reference one or more attached spreadsheets. For each selected thread, the model summarizes the communicative intent and articulates a workflow description.
We distinguish two cases. In the strongly grounded case illustrated in Figure 4 (a), both the input and the final reference artifacts for the workflow are already present as attachments in the thread (e.g., an initial ranking file and a corrected version). Strongly grounded cases well motivated this benchmark, but they are relatively rare. In the partially grounded case illustrated in Figure 4 (b), the email specifies a clear goal, but only some of the required artifacts are attached (e.g., just the updated schedule, or an intermediate report). Across both cases, human experts normalize conversational email text into task instructions and abstract away idiosyncratic details while preserving the business intent. For strongly grounded threads, annotators primarily verify that the attached files exactly implement the requested change without extra edits. For partially grounded threads, they either identify the missing artifacts from versioned spreadsheets and revise them to align with the described workflow—carefully avoiding the introduction of new changes—or create the missing artifacts themselves, which typically requires much more effort.
@@ 2.1.2 Workflow Derivation from Versioned Spreadsheets
2.1.3 Workflow Sourced from High-quality Artifacts
Finally, we curate workflows from high-quality spreadsheets and reports drawn from the Enron and EUSES corpora, various investment and securities companies, international organizations, and national governments (e.g., the World Bank and the Canadian and British governments). Domain experts decompose these finalized financial artifacts into intermediate artifacts that serve as inputs, while treating the original documents as reference outputs, and then author realistic workflow instructions. For example, a valuation model from an investment firm can be turned into a financial modeling task; a World Bank report can be used to define a data-driven summarization and visualization task; and bilingual reports from the Canadian government can be used to construct translation and consistency-checking tasks. We additionally leverage labeled samples from existing datasets: we adapt 10 financial cases from WideSearch [wong2025widesearch] into web-search-centric workflows and extend them into multi-step calculation and visualization pipelines, and we leverage 3 examples from DABStep [egg2025dabstep] to construct multi-source question answering workflows.
@@ 2.2 Dataset Characteristics
Finch comprises 172 meticulously annotated, enterprise-grade workflows that collectively capture the compositional, messy, multimodal, and collaborative nature of real finance and accounting work. Across these workflows, the corpus contains 1,710 spreadsheets together with 13 PDFs, 7 images, 3 Word documents, and additional files such as JSON, CSV, and Markdown. This mixture reflects how real analysts coordinate over heterogeneous artifacts rather than clean, single-table inputs.
@@ 2.2.1 Task Compositionality
Finch is explicitly designed around composite workflows rather than isolated tasks. As shown in Figure 5, only 37 workflows (21.5%) are single-task; the remaining 135 (78.5%) involve multiple tasks. Importantly, each “task” itself typically requires substantial multi-turn reasoning: for example, web search often entails many rounds of LLM calls to discover, filter, and verify evidence; cross-sheet retrieval requires repeated calls to read and locate key information across multiple sheets; and calculation usually spans many formulas distributed over different rows and columns.
@@ 2.2.2 Messiness
The source files in Finch are intentionally large, multi-sheet, and structurally complex. At the file level, 86.6% of workflows involve more than one file when counting both input and reference artifacts, and some workflows touch up to 14 distinct files. At the spreadsheet level, 92.4% of workflows involve multiple sheets, with an average of 8 sheets and a long tail reaching 91 sheets. As a result, systems must navigate cross-sheet dependencies, hidden logic, and scattered intermediate calculations rather than operating on a single “analysis” sheet. Moreover, most spreadsheets exhibit complex layouts that interleave text, tables, formulas, and charts, as well as intricate table structures with merged cells, nested headers, hierarchical data, blank rows and columns, and other irregularities.
Cell-level statistics further highlight the scale of the data. The median workflow covers 15K cells (157K on average for all workflows), with the largest workbook scaling to 3.7 million cells. Formula density is similarly skewed: while workflows contain an average of 21.5K formulas (the median is 212), the most complex ones have hundreds of thousands of formulas, reflecting deeply nested calculations and long dependency chains. Taken together, these properties create a challenging regime in which models must reason over large, noisy, and highly irregular spreadsheet layouts, rather than clean, rectangular tables.
@@ 2.3.2 LLM-as-Judge Evaluation
For generation tasks involving spreadsheets, the framework extracts all cell values and formulas from both the reference and the model output, and captures screenshots of every sheet, since the entire generated artifact must be verified rather than just localized edits. For QA tasks, it feeds a rich text-and-image sequence constructed from the reference answer and the model’s response, optionally augmented with relevant input files when the question requires grounding in external artifacts.
@@ 3.1.2 API-based Models
We evaluate five frontier large language models (LLMs), shown in Table 1, through API interfaces. Our evaluation pipeline builds upon the open-source SpreadsheetBench framework [ma2024spreadsheetbench]. Details of the execution paradigm and prompting strategies are deferred to the Appendix B.1. While SpreadsheetBench provides only a baseline implementation, we introduce several practical enhancements to improve its applicability to real-world tasks.
@@
SpreadsheetBench only produces simple text tables without preserving cell addresses, data types, or formulas. However, these details are essential for tasks in Finch. We extend SpreadsheetLLM [spreadsheetllm2024] encoding and introduce a semantic-rich tuple encoding that preserves full structural and semantic fidelity. Each sheet begins with its name and the corresponding data bounding box (e.g. ## Sheet: [name] (A1:Z100)). We then serialize the bounded region using a Markdown-based format. Each cell is encoded as a tuple (Address, Value, Type, Formula), where Address denotes the cell reference (e.g. A3), Type indicates the data type (T = Text, I = Integer, F = Float, D = Date, B = Boolean), and Formula records the cell formula (e.g., =SUM(A1:A10)->100).
@@ 3.2 Experimental Results
Web agents (ChatGPT 5.1 Pro vs. Claude Sonnet 4.5).
@@
Figure 8: Pass rate comparison for GPT 5.1 Pro and Claude Sonnet 4.5 across different task combinations and task types. The left chart visualizes the aggregated pass rate based on task combinations, revealing the models’ capabilities in handling multi-step workflows commonly seen in professional finance and accounting tasks. The $k$ in the right chart represents the number of tasks included in a workflow. For example, a "k=3" workflow involves three distinct tasks, and its pass rate is calculated based on the collective performance of those tasks. The right chart shows the pass rates for individual tasks performed by both models in the Finch benchmark. For workflows that involve multiple tasks (i.e., composite workflows), the correctness is counted for each individual task, as it is difficult to isolate which specific task caused the failure.
@@
Figure 2 further shows that our API-based method remains competitive despite using only a single LLM call. In automated evaluation, for example, GPT 5.1 Pro achieves 41.9%, while GPT 5.1 with our API-based agent design reaches 32.0%—a relatively small gap given the much weaker interaction budget. Our agent design narrows the gap by supplying more efficient spreadsheet representations and task-appropriate tool outputs within the single-call budget, thereby reducing brittle reasoning over complex layouts and mitigating cascading mistakes.
@@ 3.2.1 Consistency Between Human and Automated Evaluation
We adopt a lightweight multimodal model, GPT-5-mini, as the judge to validate the performance of our automated evaluation framework. As shown in Table 3, the automated evaluation largely aligns with human judgments: for GPT 5.1 Pro and Claude Sonnet 4.5, the judge agrees with human labels on 82.1% and 90.2% of workflows, respectively. The judge also achieves high recall (83.3% and 88.4%), meaning it recovers most human-labeled passes, and reasonably strong precision (73.6% and 76.0%), indicating that the majority of automatically predicted passes are also accepted by human evaluators. Overall, this suggests that automated evaluation may overestimate accuracy by several percentage points.
@@ 3.3 Error Analysis
To understand the sources of failure on Finch, we conducted a qualitative error analysis of GPT 5.1 Pro and Claude Sonnet 4.5 in both their web-agent and API configurations. For all failed workflows in our evaluation, we manually inspected the trajectories and annotated the primary cause of failure.
From a workflow-centric perspective, we identify five dominant categories of error. [^1] For this study, once we identify the first clear error in a failed workflow, we stop further analysis for that workflow. Take Claude Sonnet 4.5 web-agent as an example. Across all examined failures, 10% stem from task misunderstanding: enterprise tasks often rely on implicit context in enterprise artifacts (e.g., spreadsheets), which models frequently overlook, leading them to misinterpret what is being asked and the required deliverable. 25% are data retrieval errors, including selecting the wrong cross-sheet, cross-table, or intra-table row/column ranges. 35% arise from formula reasoning errors, such as failing to reconstruct the latent business logic encoded in formulas or deriving incorrect new formulas. 25% are due to code generation errors, where generated scripts (e.g., Python with spreadsheet APIs) are syntactically invalid or misaligned with the spreadsheet layout. The remaining 5% correspond to data rendering errors, including incorrect formatting, misconfigured charts, or flawed final reports that deviate from the requested layout or narrative—for example, creating a brand-new spreadsheet instead of modifying the original one as requested. We also compare error patterns between web-based agents and API-based setups, with details provided in Appendix C.

### v3

@@ Abstract
We introduce a finance & accounting benchmark (Finch) for evaluating AI agents on real-world, enterprise-grade professional workflows—interleaving data entry, structuring, formatting, web search, cross-file retrieval, calculation, modeling, validation, translation, visualization, and reporting. Finch is sourced from authentic enterprise workspaces at Enron (15,000 spreadsheet files and 500,000 emails from 150 employees) and other financial institutions, preserving in-the-wild messiness across multimodal artifacts (text, tables, formulas, charts, code, and images) and spanning diverse domains such as budgeting, trading, and asset management.
@@
We conduct both human and automated evaluations of frontier AI systems including GPT 5.1, Claude Sonnet 4.5, Gemini 3 Pro, Grok 4, and Qwen 3 Max. GPT 5.1 Pro spends 16.8 minutes per workflow yet passes only 38.4% of workflows, while Claude Sonnet 4.5 passes just 25.0%. Comprehensive case studies further surface the challenges that real-world enterprise workflows pose for AI agents.
@@ 1 Introduction
However, real-world F&A work is inherently messy, with substantial contextual complexity: artifacts are interconnected across heterogeneous spreadsheets, PDFs, and other artifacts, evolving through multiple versions with collaborative edits [klimt2004enron]; spreadsheets contain large, complex structures [spreadsheetllm2024] with cross-sheet references, intricate layouts, inconsistent formatting, cryptic terms, erroneous formulas, and multimodal artifacts such as charts, images, and code. It is also long-horizon [gdpval2025]: workflows demand multi-step reasoning spanning data entry, editing, retrieval, calculation, modeling, validation, reporting, and more.
@@
  • In-the-wild enterprise sourcing: Finch is built around authentic enterprise spreadsheets, emails, and PDFs from real-world enterprise workspaces—primarily Enron [enrondata_pst] (about 15,000 spreadsheet files and 500,000 emails from 150 executives and employees) and EUSES [fisher2005euses] (about 450 financial spreadsheet files from various sources), along with securities and asset management firms, global organizations such as World Bank [worldbank_idr_2024], and Canadian and British governments [financecanada_fiscalref_2025, hmt_pesa_2023]. Documents are large, cross-referenced, and messy—containing rich multimodal artifacts such as tables, formulas, charts, pivots, and images.
  • Rigorous construction process: We propose a novel workflow construction pipeline grounded in real collaborative context of emails and versioned artifacts. We induce workflows from enterprise email threads and attachments, where collaborators naturally describe, discuss, and track workflows as part of their daily work. We also propose an LLM-assisted, expert-verified method to derive workflows by analyzing changes across versioned spreadsheets, surfacing the underlying goals that drive professionals’ work. Annotators must reason over large multi-sheet workbooks, and subtle version deltas to infer what the original analyst was trying to achieve, making the annotation process substantially more difficult than curating QA pairs over isolated tables.
@@
Figure 3: Illustration of an end-to-end predictive modeling workflow typically performed by financial analysts. The workflow involves multiple steps, including web search, data import, cross-sheet and cross-file retrieval, calculation and financial modeling, visualization, and report generation. More illustrative examples for data characteristics in Finch are presented in Appendix E.
@@ 2.1 Dataset Construction
All annotated workflows from these different channels are consolidated into a unified schema with consistent fields (NL instruction, input files, reference outputs), and each workflow is tagged with task types (e.g., data entry/import, structuring, validation) and business types (e.g., planning and budgeting, pricing and valuation, operations, asset management). Note that the reference outputs may include both file-based reference answers (for most generation/editing cases) and textual reference answers (for a small number of QA and summary/visualization cases).
@@ 2.1.1 Workflow from Enterprise Email Threads
We first mine real-world enterprise email threads to surface workflows. Starting from the Enron Email Corpus, we prompt GPT-5 to identify collaborative messages that (i) explicitly state a business goal (e.g., “update the RAC rankings” or “revise the 2002 allocations”) and (ii) reference one or more attached spreadsheets. For each selected thread, the model summarizes the communicative intent and articulates a workflow description.
In the strongly grounded case illustrated in Figure 4 (a), both the input and the final reference artifacts for the workflow are already present as attachments in the thread (e.g., an initial ranking file and a corrected version). Strongly grounded cases were a primary motivation of this benchmark, but they are relatively rare. In the partially grounded case illustrated in Figure 4 (b), the email specifies a clear goal, but only some of the required artifacts are attached (e.g., just the updated schedule, or an intermediate report). Across both cases, human experts normalize conversational email text and LLM-drafted descriptions into workflow instructions and abstract away idiosyncratic details while preserving the business intent. For strongly grounded threads, annotators primarily verify that the attached files exactly implement the requested change without extra edits. For partially grounded threads, they either identify the missing artifacts from attached spreadsheets and revise them to align with the described workflow—carefully avoiding the introduction of new changes—or create the missing artifacts themselves, which typically requires much more effort.
@@ 2.1.2 Workflow Derivation from Versioned Spreadsheets
2.1.3 Workflow Sourced from Final Deliverable Spreadsheets and Reports
Finally, we curate workflows from high-quality spreadsheets and reports drawn from the Enron and EUSES corpora, various investment and securities companies, international organizations, and national governments (e.g., the World Bank and the Canadian and British governments). Domain experts author realistic workflow instructions and construct input and reference files based on final deliverable artifacts. For example, a valuation model from an investment firm can be turned into a financial modeling task; a World Bank report can be used to define a data-driven summarization and visualization task; and bilingual reports from the Canadian government can be used to construct translation and consistency-checking tasks. We additionally leverage labeled samples from existing datasets: we adapt 10 financial cases from WideSearch [wong2025widesearch] into web-search-centric workflows and extend them into multi-step calculation and visualization pipelines, and we leverage 3 examples from DABStep [egg2025dabstep] to construct multi-source question answering workflows.
@@ 2.2 Dataset Characteristics
Finch comprises 172 meticulously annotated, enterprise-grade workflows that collectively capture the compositional, messy, multimodal, and collaborative nature of real finance and accounting work. Across these workflows, the corpus contains 1,710 spreadsheets (956 distinct spreadsheets) together with 17 PDFs, 12 images, 3 Word documents, and additional files such as JSON, CSV, and Markdown. This mixture reflects how real analysts coordinate over heterogeneous artifacts rather than clean, single-table inputs.
@@ 2.2.1 Task Compositionality
Finch is explicitly designed around composite workflows rather than isolated tasks. As shown in Figure 5, only 37 workflows (21.5%) are single-task; the remaining 135 (78.5%) involve multiple tasks. Importantly, each “task” itself typically requires substantial multi-turn reasoning: for example, web search often entails many rounds of LLM calls to discover, filter, and verify evidence; cross-sheet retrieval requires iterative calls to read and locate key information across multiple sheets; and calculation usually spans many formulas distributed over different rows and columns.
@@ 2.2.2 Messiness
The source files in Finch are intentionally large, multi-sheet, and structurally complex. At the file level, 86.6% of workflows involve more than one file when counting both input and reference artifacts, and a workflow touches up to 14 distinct files. At the spreadsheet level, 92.4% of workflows involve multiple input and reference sheets, with an average of 8 sheets and a long tail reaching 91 sheets. As a result, systems must navigate cross-sheet dependencies, hidden logic, and scattered intermediate calculations rather than operating on a single “analysis” sheet. Moreover, most spreadsheets exhibit complex layouts that interleave text, numerical values, formulas, and charts, as well as intricate single- or multi- table structures with nested headers, hierarchical data, merged cells, blank rows and columns, and other irregularities.
Cell-level statistics further highlight the scale of the data. The median workflow covers 15K cells (157K on average for all workflows), with the largest one scaling to 3.7 million cells. Formula density is similarly skewed: while workflows contain an average of 21.5K formulas (the median is 212), reflecting deeply nested calculations and long dependency chains. Taken together, these properties create a challenging regime in which models must reason over large, noisy, and highly irregular spreadsheet layouts, rather than clean, rectangular tables.
@@ 2.3.2 LLM-as-Judge Evaluation
For generation tasks involving spreadsheets, the framework extracts all cell values and formulas from both the reference and the model output, and captures screenshots of every sheet, since the entire generated artifact must be verified rather than just localized edits. For QA tasks, it feeds the reference answer and the model’s response, optionally augmented with relevant input artifacts when the question requires grounding in input artifacts.
@@ 3.1.2 API-based Models
We evaluate five frontier LLMs via API interfaces (Table 1). We adopt SpreadsheetBench [ma2024spreadsheetbench] as the baseline framework because it provides a principled code-generation paradigm for spreadsheet-centric tasks, treating executable code as the model’s action space and enabling direct manipulation of spreadsheets through standard libraries. This design naturally aligns with Finch, where workflows require complex spreadsheet operations, formula reasoning, and cross-sheet dependencies that cannot be reliably handled by text-only outputs.
While SpreadsheetBench was originally designed for relatively small and clean spreadsheets, we extend it with richer spreadsheet encodings, multimodal input support, and stricter execution and evaluation protocols, allowing it to scale to the large, messy, and long-horizon enterprise workflows in Finch.
@@
SpreadsheetBench produces text tables without preserving cell addresses, data types, or formulas. However, these details are essential for tasks in Finch. We extend SpreadsheetLLM [spreadsheetllm2024] encoding and introduce a semantic-rich tuple encoding that preserves full structural and semantic fidelity. Each sheet begins with its name and the corresponding data bounding box (e.g. ## Sheet: [name] (A1:Z100)). We then serialize the bounded region using a Markdown-based format. Each cell is encoded as a tuple (Address, Value, Type, Formula), where Address denotes the cell reference (e.g. A3), Type indicates the data type (T = Text, I = Integer, F = Float, D = Date, B = Boolean), and Formula records the cell formula (e.g., =SUM(A1:A10)->100).
@@ 3.2 Experimental Results
Product-side agents (ChatGPT 5.1 Pro vs. Claude Sonnet 4.5).
@@
Figure 8: Pass rate comparison for GPT 5.1 Pro and Claude Sonnet 4.5 across different task combinations and task types. The left chart visualizes the aggregated pass rate based on task combinations, revealing the models’ capabilities in handling multi-step workflows commonly seen in professional finance and accounting tasks. The $k$ in the right chart represents the number of tasks included in a workflow. For example, a "k=3" workflow involves three distinct tasks, and its pass rate is calculated based on the collective performance of those tasks. The right chart shows the pass rates for individual tasks performed by both models in the Finch benchmark. For workflows that contain multiple tasks, a task is counted as correct only if the entire workflow is completed successfully. If a workflow fails, all tasks within that workflow are counted as incorrect.
@@
Figure 2 further shows that the adapted API-based baselines are generally weaker than official product-side agents. Under automated evaluation, for example, GPT 5.1 Pro achieves a pass rate of 41.9%, whereas GPT 5.1 with our API-based agent design reaches 32.0%. A key limitation of the API-based baselines is that they rely on a single LLM call, which precludes iterative interaction, execution feedback, and self-correction—an important direction for future work in designing F&A enterprise-grade agent frameworks. Despite this constraint, our agent design narrows the performance gap by employing more efficient spreadsheet encodings and task-appropriate tool outputs within the single-call budget.
@@ 3.2.1 Consistency Between Human and Automated Evaluation
We adopt a lightweight multimodal model, GPT-5-mini, as the judge for automated evaluation framework. As shown in Table 3, the automated evaluation largely aligns with human judgments: for GPT 5.1 Pro and Claude Sonnet 4.5, the judge agrees with human labels on 82.1% and 90.2% of workflows, respectively. The judge also achieves high recall (83.3% and 88.4%), meaning it recovers most human-labeled passes, and reasonably strong precision (73.6% and 76.0%), indicating that the majority of automatically predicted passes are also accepted by human evaluators. Overall, this suggests that automated evaluation may overestimate accuracy by several percentage points.
@@ 3.3 Error Analysis
To understand the sources of failure on Finch, we conducted a qualitative error analysis of GPT 5.1 Pro and Claude Sonnet 4.5 in both their product-agent and API configurations. For all failed workflows in our evaluation, we manually inspected the trajectories and annotated the primary cause of failure.
From a workflow-centric perspective, we identify five dominant categories of error. [^1] For this study, once we identify the first clear error in a failed workflow, we stop further analysis for that workflow. Take Claude Sonnet 4.5 product-agent as an example. Across all examined failures, 10% stem from task misunderstanding: enterprise tasks often rely on implicit context in enterprise artifacts (e.g., spreadsheets), which models frequently overlook, leading them to misinterpret what is being asked and the required deliverable. 25% are data retrieval errors, including selecting the wrong cross-sheet, cross-table, or intra-table row/column ranges. 35% arise from formula reasoning errors, such as failing to reconstruct the latent business logic encoded in formulas or deriving incorrect new formulas. 25% are due to code generation errors, where generated scripts (e.g., Python with spreadsheet APIs) are syntactically invalid or misaligned with the spreadsheet layout. The remaining 5% correspond to data rendering errors, including incorrect formatting, misconfigured charts, or flawed final reports that deviate from the requested layout or narrative—for example, creating a brand-new spreadsheet instead of modifying the original one as requested. We also compare error patterns between web-based agents and API-based setups, with details provided in Appendix C.